# Intergroup violence in bursts or fizzles

### v10

@@ 3 Model
The critical threshold of turmoil increases with network size at a decreasing rate [10], but it also increases with the proportion of unconditional defectors. At $T_{c}$, simulations point out that cooperation starts in small clusters of conditional cooperators, which we will also test empirically. The onset in small groups is puzzling because larger groups have a better chance to win at lower individual costs. Yet if in a small (sub)group, someone starts fighting, he (and rarely she) accounts for a relatively sizeable proportion of his neighbors’ social contacts, and more readily wins them over to fight than in a large group where he would comprise a small minority.
Fighting ends when exhaustion sets in, a winner stands out (opponents have fled or have been wrestled to the ground), or others intervene.
@@ 4 Data and Methods
In studying violence, lab experiments lack the turmoil and emotional intensity of violent confrontations due to their obligation to meet ethical standards. Field studies, in contrast, cannot be based on a random sample of participants or groups, yet they are invaluable for realistic view on violence [4]. We obtained 42 videos from websites such as YouTube, LiveLeak, and WorldStarHipHop using search terms with the English keywords “brawl,” “street fight,” and “assault.” This sample is not random with respect to violence, but it is random with respect to temporal unfolding and (sub)group size. Of these clips, 36 are from English-speaking countries (mainly the US and the UK, with one from Canada and one from India); five of the remaining clips are from the Netherlands, and one is from Colombia. We did not observe differences in relevant behavior related to the location of the recording. To keep distracting factors away from our analysis, we excluded clips with professional fighters, long range weapons, protective clothing, a referee, ambush attacks, or youths in a school yard. Most of our videos are recorded on phones by bystanders and are left-truncated; in all likelihood, there would have already been some turmoil that motivated bystanders to start filming. The shortest lasted 30 sec. and the longest was nearly 5 minutes (mean 101 sec.; s.d. 59 sec.). Out of a potential 2 x 42 groups, where the opponents in one analysis become the focal group in the next, 25 groups attacked a single individual rather than a group. Because a lone individual is unable to act collectively, this leaves 59 groups to examine; see the Supplementary information for all case studies and elaborate examples of interpreting and coding. One group had 14 members, but all other groups were small, $102\leq n<$ 10 (mean 3.6). They were simulated as cliques wherein everyone could sense one another unless there were obstacles or de-escalators obstructing contact.
@@
We coded violence when force was used against another’s body (punching, slapping, kicking, hitting, stomping) and/or when another person’s body was forcefully moved. People may defect (not use violence) for all kinds of reasons and due to various causes. To contrast everyone not fighting from conditional cooperators, we coded them as unconditional defectors if others were fighting while they were not. Among them we include de-escalators, as well as group members who were prevented to fight by de-escalators.
@@
Given the distraction caused by turmoil, it is not feasible for all group members to react within 1 second to a first cooperator (as they might in a well-organized sports team), whereas 3 seconds is too long for the notion of burst to apply to our small groups. Hence we defined a burst as an outbreak of violence by at least half of the group (Fig. 2), or both individuals in a dyad, if they started fighting less than 2 seconds after the first, with a 5% margin. [^7] The requirement that $N_{C}\geq 0.5n$ is based on simulations of small networks just above $p_{c}$, but under this condition in large networks (e.g., $n=1000$), $N_{C}<0.5n$.
@@ 4.1 Ethics
The use of videos for research purposes poses distinct ethical challenges, largely due to the nonanonymous content of the videos. However, ethical guidelines for digital spaces tend to be less restrictive [48], with the consent of the participants being less stringent for data acquired from the public domain, including the internet. While our video corpus is open for use and inspection by other researchers upon request, we require that they take the same measures to ensure the anonymity of the persons portrayed as we did.
@@ 5 Results
We confirm the finding of an earlier study [19] that fighting tends to start in small groups or in small subgroups of larger groups. As predicted [10], we found that small (sub)groups burst into action at lower levels of turmoil than larger groups, and lower levels are of course reached earlier. The local emergence (versus central coordination) of collective action in small groups has also been observed in protests [71]. In our data, bursts developed in 13 (37%) of the 35 smallest groups (dyads and triads) and in 10 (42%) of the 24 larger groups. In bursts, the correlation between focal group size and opponents’ turmoil is 0.53. The size-turmoil relationship is slightly disturbed by dyads more often facing a larger opponent group and being therefore less likely to fight collectively than triads, which in turn are more robust in general [66]. [^8] The predicted relation between group size and turmoil is $T\approx n^{1/2}$ [10], but due to the small data and unexplained variance therein, a curve does not fit better than a straight line; hence the reported correlation. If there was a clear winner, it was always the larger group, with only one exception, in line with earlier models of warfare controlling for weapons [38].
@@
The proportions of unconditional defectors in groups with bursts (mean = 0.19; s.d. = 0.21) and groups without bursts (mean = 0.49; s.d. = 0.27) are box-plotted in Figure 4 (Welch test $t=4.796;P=6.411$ x $10^{-6}$; df = 54.76; the ROC-curve has AUC = 0.760). The predicted critical threshold (vertical line in the figure) separates the two boxplots, but one might question how robust this result is, given some noise in coding. There is a risk of missing a violent act (it can be extremely fast) or mislabeling a fast movement as violence, thus mixing up fighters and defectors, and a risk of misassigning individuals to groups. Even if we suppose there is a 10% chance of each of these errors (which corresponds to Krippendorff’s $\alpha=0.75$; see inter-coder reliability in the Supplementary information), the result is hardly weakened (Welch test $t=3.879;P=0.0025$, averaged over 1000 simulation runs). Hence the distinction between bursts and fizzles appears to be robust.
@@
Figure 4: The proportion of unconditional defectors and the predicted critical level (vertical line), which for separates the boxplots of bursts (1) and fizzles (0). The data points are our 59 groups.
@@
The empirical distinction between bursts and fizzles is not perfect, and Figure 4 shows that 15 groups out of 59 (25%) are categorized incorrectly. In 11 of these 15 groups, the time between the first violent act and last participant joining in was above our two-second interval, hence we did not count these collective actions as bursts. In simulations, lags between first instances of violence and the moment when at least half of the group participates frequently occur when turmoil lingers at its critical level (Fig. 5), versus steadily increasing turmoil that yields a clean burst (Fig. 2). Lingering turmoil seems to have occurred in 4 cases, but we would need a better scale of turmoil than the cumulative scale we currently use to be sure. Far more important are local circumstances beyond the model. When re-examining the time lags beyond two seconds, we noticed various causes of delay: (1) group members were in a car and it took time to get out; (2) a focal group member stood at a distance from the fight on a slippery floor next to a pool; (3) a focal group member was beaten and seemed intoxicated by alcohol, hence responded slowly; (4) a focal group member was first pushed over and it took him time to get on his feet again; (5) opponents moved around so fast that it took time to hit one; (6) an opponent fell into thick bushes such that only one focal group member at a time could get to him; (7) group members were sitting in a train and needed time to get on their feet; (8) group members were constrained in a small room where they searched for space to lash out; and, (9) group members were hindered by the mess of two opponents fighting on a collapsed table full of food. One can hardly blame the model for not incorporating these local circumstances that caused delays, and perhaps our two seconds limit should be applied with more lenience in such cases. Finally, two delays were caused by ambiguous behavior: (10) someone tried to de-escalate, was hit, and then attacked, and (11) someone tried to pull back his fighting friend (de-escalation) but simultaneously attacked.
@@
Figure 5: Simulation of Equation 1 at a the critical level of turmoil, $T_{c}\approx 0.11$, over a range of Monte Carlo steps as proxies for time points. The vertical axis shows the proportion of cooperators ($N_{C}/n$) in a fully connected network ($n=5$) with one unconditional defector. Note an extended period of multiple short spikes of violence before all conditional cooperators join in. Similar lags occur without unconditional defectors.
@@
In the 4 out of 15 remaining groups, non-bursts were predicted whereas bursts occurred. Although for parsimony, we applied the same $S_{0}$ (and $\Delta$) for all conditional cooperators, some combatants may actually have a higher value than the average, $S_{0,i}>\overline{S_{0}}$, which could explain these cases.
@@ 6 Discussion and Conclusion
The simple Ising model is a century old [37] and has been applied to a wide range of problems [70, 43], to which we add the dilemma of collective action. It explains cooperation parsimoniously, based on agitating stimuli without recourse to strong rationality, initiative takers, norms, feedback through selective incentives, or reliable information passing through the network, resulting in reputations. The Ising model elucidates the temporal unfolding of violence by predicting a critical threshold of unconditional defectors that distinguishes a burst of collective action from a fizzle, which is largely supported by the data. The model also explains why violent groups are often small or are small subgroups of larger groups despite greater risk. Small (sub)groups have a lower critical threshold of turmoil, and in a confrontation with opponents, lower levels are reached earlier. We also investigated whether synchronous action precedes violence, but we found that synchronization precedes both collective and solitary violence, and cannot predict either of these outcomes. However, synchronization may still be important to increase solidarity [27, 22].
This study has several limitations. Because we did not select videos without violence, we cannot be certain that violence is caused by turmoil. Moreover, developing a proper scale of turmoil or agitation for field studies is exceedingly difficult, due to variations across domains (e.g., street violence versus grievances about politics), and different effect sizes of different stimuli (e.g., killed family members versus a push that might be quickly forgotten). On top of these challenges, our measurements underestimated turmoil because the videos are left-truncated, and our data depend on camera angle and vision width, inevitably excluding instances of turmoil. Measuring turmoil or agitation is only possible in exceptional situations, for example when protest movements produce it by themselves through accelerated posting of online messages. In these particular cases, the inter-posting time intervals follow a certain pattern (Moore’s law), which makes it possible to predict $T_{c}$ ahead of large street demonstrations [33]. In lab experiments, it is unlikely that the patterns of bursts and fizzles would have been discovered, but now they have been pointed out in the field, ethically responsible intergroup conflicts might be examined experimentally, benefiting from controlled circumstances and unambiguous measurement of agitation. Then, causality of turmoil and of the critical level of unconditional defectors might be established or refuted.
@@
In this first empirical application, we showed that the Ising model can explain the unfolding of violence, and in all likelihood, more discoveries lay ahead. Extensions to norms (as external fields) and noise in actors’ information about others’ behavior (i.e., reputations) have been explored in simulations [9]. In all likelihood, it can be applied to protests [29] and revolts, which break out more often if (rumors say that) a government or its police are weakened [67, 74], analogous to vulnerable individuals in street fights. The model also seems applicable to bystanders collectively helping victims under uncertainty [58], and to lynchings [6]. For example, Nussio [52] argued that lynchings in Mexico can be explained through solidarity combined with peer pressure in a network transmitting reputations meshed with local norms. This argument is perfectly consistent with an expanded Ising model [9], but it lacks the agitation necessary for cooperation to start without centralized organization, as it did. We would argue that the agitation was due to the violation of social norms, of which the lynching victims were accused (e.g., child theft), augmented by the gossip transmitted through the network. In other countries, putative norm violations that agitated crowds and lead to lynchings were rape [12] and blasphemy [3, 60].
Taking a broader perspective on the role of random noise, multiple studies have found that it can solve coordination problems, such as collective responses to opponents’ threats, by shaking a group loose from its suboptimal (e.g., non-cooperative) state [65], but it can also disturb an optimal state. In the Ising model, both can happen, depending on the amount of turmoil: when in a group, turmoil approaches $T_{c}$ from below, random noise in decision making facilitates cooperation, but at high levels, too much noise entails confusion and error.
@@ 7 Data and compilation of data set
The shortest clip lasted 30 sec. and the longest nearly 5 minutes (mean 101 sec.; s.d. 59 sec.). Out of a potential 2 x 42 groups, where the opponents in one analysis become the focal group in the next, 25 groups fought with a single individual rather than a group, which leaves $\mathcal{N}$=59 groups to examine. Most groups were small, $102\leq n<$ 10 (mean 3.6), but one had 14 members (of which 6 became violent). The smaller ones were simulated as fully connected networks wherein everyone could see one another unless there were obstacles or deescalators obstructing visual contact. Obstruction was simulated by removing $m$ ties.
@@ 8 Coding
[…]

### v12

@@ 3 Model
The critical threshold of turmoil increases with network size at a decreasing rate [10], but it also increases with the proportion of unconditional defectors. At $T_{c}$, simulations point out that cooperation starts in small clusters of conditional cooperators, which we will also test empirically. The onset in small groups is puzzling because larger groups have a better chance to win at lower individual costs. Yet if in a small (sub)group, someone starts fighting, he (and rarely she) accounts for a relatively sizeable proportion of his neighbors’ social contacts, and more readily wins them over to fight than in a large group where he would comprise a small minority. Fighting ends when exhaustion sets in, one party dominates (i.e., opponents flee or are wrestled to the ground), or others intervene.
@@ 4 Data and Methods
In studying violence, lab experiments lack the turmoil and emotional intensity of violent confrontations due to their obligation to meet ethical standards. Empirical field studies, in contrast, cannot be based on a random sample of participants or groups, yet they are invaluable to obtain a realistic view of violence [4]. We obtained 42 videos from websites such as YouTube, LiveLeak, and WorldStarHipHop using search terms with the English keywords “brawl,” “street fight,” and “assault.” This sample is not random with respect to violence, but it is random with respect to temporal unfolding and (sub)group size. Of these clips, 36 are from English-speaking countries (mainly the US and the UK, with one from Canada and one from India); five of the remaining clips are from the Netherlands, and one is from Colombia. We did not observe differences in relevant behavior related to the location of the recording. To keep distracting factors away from our analysis, we excluded clips with professional fighters, long range weapons, protective clothing, a referee, ambush attacks, or youths in a school yard. Most of our videos are recorded on phones by bystanders and are left-truncated; in all likelihood, there would have already been some turmoil that motivated bystanders to start filming. The shortest lasted 30 sec. and the longest was nearly 5 minutes (mean 101 sec.; s.d. 59 sec.). Out of a potential 2 x 42 groups, where the opponents in one analysis become the focal group in the next, 25 groups attacked a single individual rather than a group. Because a lone individual is unable to act collectively, this leaves 59 groups to examine; see the Supplementary information for all case studies and elaborate examples of interpreting and coding. One group had 14 members, but all other groups were small, $2\leq n<10$ (mean 3.6). They were simulated as cliques wherein everyone could sense one another unless there were obstacles or de-escalators obstructing contact.
@@
We coded violence when force was used against another’s body (punching, slapping, kicking, hitting, stomping) and/or when another person’s body was forcefully moved. People may defect (not use violence) for all kinds of reasons and due to various causes. To contrast everyone not fighting from conditional cooperators, we coded them as unconditional defectors if others were fighting while they were not. Among them we include de-escalators, as well as group members who were prevented to fight by de-escalators (who in turn might belong to either group or none).
@@
Given the distraction caused by turmoil, it is not feasible for all group members to react within 1 second to a group member who initiates violence (as they might in a well-organized sports team), whereas 3 seconds is too long for the notion of burst to apply to our small groups. Hence we defined a burst as an outbreak of violence by at least half of the group (Fig. 2), or both individuals in a dyad, if they started fighting less than 2 seconds after the first, with a 5% margin. [^7] The requirement that $N_{C}\geq 0.5n$ is based on simulations of small networks just above $p_{c}$, at this proportion in large networks (e.g., $n=1000$), $N_{C}<0.5n$.
@@ 4.1 Ethics
The use of videos for research purposes poses distinct ethical challenges, largely due to the non-anonymous content of the videos. However, ethical guidelines for digital spaces tend to be less restrictive [52], with the consent of the participants being less stringent for data acquired from the public domain, including the internet. While our video corpus is open for use and inspection by other researchers upon request, we require that they take the same measures to ensure the anonymity of the persons portrayed as we did.
@@ 5 Results
We confirm the finding of an earlier study [20] that fighting tends to start in small groups or in small subgroups of larger groups. As predicted [10], we found that small (sub)groups start fighting at lower levels of turmoil than larger groups, and lower levels are of course reached earlier. The local emergence (versus central coordination) of collective action in small groups has also been observed in protests [76]. The size-turmoil relationship is slightly disturbed by dyads more often facing a larger opponent group and being therefore less likely to fight collectively than triads. If there was one party that dominated, it was always the larger group, with only one exception, in line with earlier models of warfare when controlling for weapons [41].
In our data, bursts developed in 13 (37%) of the 35 smallest groups (dyads and triads) and in 11 (46%) of the 24 larger groups. In bursts, the correlation [^8] The predicted relation is $T_{c}\approx n^{1/2}$ [10], but due to the small data and unexplained variance therein, a curve does not fit better than a straight line; hence, the reported correlation. between focal group size and opponents’ turmoil is 0.53.
@@
The proportions of unconditional defectors in groups with bursts (mean = 0.19; s.d. = 0.21) and groups without bursts (mean = 0.49; s.d. = 0.26) are box-plotted in Figure 4 (Welch test $t=4.925;P=4.119$ x $10^{-6}$; df = 54.41; the ROC-curve has AUC = 0.774). The predicted critical threshold (vertical line in the figure) separates the two boxplots, but one might question how robust this result is, given some error in coding. There is a risk of missing a violent act (that can be executed extremely rapidly) or mislabeling a fast movement as violence, thus mixing up fighters and defectors, and a risk of mis-assigning individuals to groups. Even if we suppose there is a 10% chance of each of these errors (which corresponds to Krippendorff’s $\alpha=0.75$; see inter-coder reliability in the Supplementary information), the result is barely weakened (Welch test $t=4.010;P=0.002$, averaged over 1000 simulation runs). Hence, the distinction between bursts and non-bursts is fairly robust.
@@
Figure 4: The proportion of unconditional defectors, and the predicted critical level (vertical line) that separates the boxplots of bursts (1) and non-bursts (0). The data points are our 59 groups.
@@
The empirical distinction between bursts and non-bursts is not perfect, and Figure 4 shows that 14 groups out of 59 (24%) are categorized incorrectly. Most of the errors are due to local circumstances, not a flawed Ising model. In 11 of these 14 groups, the time between the first violent act and last participant joining in was above our two-second interval. Hence, we did not count these collective actions as bursts even though the level of cooperation was higher than expected in non-bursts (Fig 2). One reason for the slow starts could be turmoil lingering at its critical level. In simulations, this lingering causes lags between first instances of violence and the moment when at least half of the group participates (Fig. 5), versus steadily increasing turmoil that yields a proper burst (Fig. 2). However, we would need a better scale of turmoil than the cumulative scale we currently have to assess lingering turmoil. Moreover, in our 11 slow start cases, local circumstances appear to have overridden the Ising dynamics. When re-examining the cases with time lags beyond two seconds, we noticed various causes of delay: (1) group members were in a car and it took time to get out; (2) a focal group member stood at a distance from the fight on a slippery floor next to a pool; (3) a focal group member was beaten and seemed intoxicated by alcohol, hence responded slowly; (4) a focal group member was first pushed over and it took him time to get on his feet again; (5) opponents moved around so fast that it took time to hit one; (6) an opponent fell into thick bushes such that only one focal group member at a time could get to him; (7) group members were sitting in a train and needed time to get on their feet; (8) group members were constrained in a small room where they searched for space to lash out; and, (9) group members were hindered by the mess of two opponents fighting on a collapsed table full of food. One can hardly blame the model for not incorporating these local circumstances that caused delays, and perhaps our two seconds limit should be applied with more lenience in such cases. Finally, two delays were caused by ambiguous behavior: (10) someone tried to de-escalate, was hit, and then attacked, and (11) someone tried to pull back his fighting friend (de-escalation) but almost simultaneously attacked.
@@
Figure 5: Simulation of Equation 1 at approximately the critical level of turmoil, $T_{c}\approx 0.11$, over a range of Monte Carlo steps as proxies for time points. The vertical axis shows the proportion of cooperators ($N_{C}/n$) in a fully connected network ($n=5$) with one unconditional defector (without the unconditional defector, $T_{c}$ is smaller). Note an extended period of multiple short spikes of solitary (and once collective) action before all conditional cooperators join in. Similar lags with spikes occur without unconditional defectors.
@@
In four cases, non-bursts were predicted whereas bursts occurred. Although for parsimony, we applied the same $S_{0}$ (and $\Delta$) for all conditional cooperators, some combatants may actually have a higher value than the average, $S_{0,i}>\overline{S_{0}}$, which could explain these cases.
@@ 6 Discussion and Conclusion
The simple Ising model is a century old [13] and has been applied to a wide range of problems [75, 46], to which we add the dilemma of collective action. It explains cooperation parsimoniously, based on agitating stimuli without recourse to strong rationality, initiative takers, norms, feedback through selective incentives, or reliable information passing through the network, resulting in reputations. The Ising model elucidates the temporal unfolding of violence by predicting a critical threshold of unconditional defectors that distinguishes a burst of collective action from a slow start. This pattern is largely supported by the data. The model also explains why violent groups are often small or are small subgroups of larger groups despite greater risk. Small (sub)groups have a lower critical threshold of turmoil, and in a confrontation with opponents, lower levels are reached earlier. We also investigated whether synchronous action precedes violence, but we found that synchronization precedes both collective and solitary violence, and cannot predict either of these outcomes. However, synchronization may still be important to increase solidarity [30, 24].
This study has several limitations. Because we did not select videos without violence, we cannot be certain that violence is caused by turmoil. Moreover, developing a proper scale of turmoil or agitation for field studies is exceedingly difficult, due to variations across domains (e.g., street violence versus grievances about politics), and different effect sizes of different stimuli (e.g., killed family members versus a push that might be quickly forgotten). On top of these challenges, our measurements underestimated turmoil because the videos are left-truncated, and our data depend on camera angle and vision width, inevitably excluding instances of turmoil. Measuring turmoil or agitation is only possible in exceptional situations, for example when protest movements produce it by themselves through accelerated posting of online messages. In these particular cases, the inter-posting time intervals follow a certain pattern (Moore’s law), which makes it possible to predict $T_{c}$ ahead of large street demonstrations [36]. In lab experiments, ethically responsible intergroup conflicts might be examined experimentally, benefiting from controlled circumstances and unambiguous measurement of agitation. Then, causality of turmoil and of the critical level of unconditional defectors might be established or refuted.
@@
In this first empirical application, we showed that the Ising model can explain the unfolding of violence, and in all likelihood, more discoveries lay ahead. Extensions to norms (as external field $hN_{m}$, with $N_{m}$ norm enforcers) and noise in actors’ information about others’ behavior (i.e., reputations) have been explored in simulations, but in this empirical study, we had no data about norms and reputations, and left them out of our model. In all likelihood, it can be applied to protests [32] and revolts, which break out more often if (rumors say that) a government or its police are weakened [72, 79], analogous to vulnerable individuals in street fights. In protests, where durations are usually counted in days instead of seconds (as our data), it might be easier to distinguish bursts from slow starts than in our small groups, where the distinction sometimes looked gradual. The model also seems applicable to bystanders collectively helping victims under uncertainty [63], and to lynchings [6]. For example, Nussio [57] argued that lynchings in Mexico can be explained through solidarity combined with peer pressure in a network transmitting reputations meshed with local norms. This argument is perfectly consistent with an expanded Ising model, but it lacks the agitation necessary for cooperation to start without centralized organization, as it did empirically. We would argue that the agitation was due to the violation of social norms, of which the lynching victims were accused (e.g., child theft), augmented by the gossip transmitted through the network. In other countries, putative norm violations that agitated crowds and lead to lynchings were rape [12] and blasphemy [3, 65]. The Ising model might even be applied to other species, for example herd bulls defending group members against attacking lions, or quorum sensing bacteria [22].
Taking a broader evolutionary perspective on the role of random noise, multiple studies have found that it can solve coordination problems, such as collective responses to opponents’ threats, by shaking a group loose from its suboptimal (e.g., non-cooperative) state [70], but it can also disturb an optimal state. In the Ising model, both can happen, depending on the amount of turmoil: when in a group, turmoil approaches $T_{c}$ from below, random noise in decision making facilitates cooperation, but at high levels, too much noise entails confusion and error.
@@ 7 Data and compilation of data set
The shortest clip lasted 30 sec. and the longest nearly 5 minutes (mean 101 sec.; s.d. 59 sec.). Out of a potential 2 x 42 groups, where the opponents in one analysis become the focal group in the next, 25 groups fought with a single individual rather than a group, which leaves $\mathcal{N}$=59 groups to examine. Most groups were small, $2\leq n<10$ (mean 3.6), but one had 14 members (of which 6 became violent). The smaller ones were simulated as fully connected networks wherein everyone could see one another unless there were obstacles or deescalators obstructing visual contact. Obstruction was simulated by removing $m$ ties.
@@ 8 Coding
To identify whether individuals belonged to a focal, opponent, or third-party group, we observed whether they remained close (at touching distance) to each other as they moved through space, whether they called each other by their names, whether they touched each other (for instance, by tapping shoulders), or whether they shared a vehicle or other object (for instance, a bag or skateboard). One of the authors and two assistants discussed each case in which we observed one or more of these behaviors to determine whether the actors were part of the same group. To assess the reliability of our coding, we asked three students who were unaware of our theoretical expectations and preliminary results to code group memberships, maximum number of participants involved in violence, and the occurrence of bursts in 37 videos, containing 54 groups and 20 solitary individuals (according to our own coding). Notice that just like biologists observing primates, much experience is required to become an accurate observer, which these students did not have. Krippendorff’s alphas for group membership, fighters, and bursts were .87, .68, and .75 respectively, or .77 on average; in the main text we reported that an error percentage of 10% results in .75. We also calculated whether the distinction between bursts and fizzles in students’ coding followed the predicted critical proportion of unconditional defectors. We arrived at the number of unconditional defectors (which the students did not code) by deducing the maximum number of participants in violence from the group size. Out of 49 groups (5 cases missing or coded as individual rather than group, as we did) students identified 29 bursts and 20 fizzles (we required that at least 2 out of 3 students agreed on the observation of a burst to denote it as such). The proportion of defectors was .18 (sd =.03=.03) in bursts and .33 (sd =.04=.04) in fizzles (Welch test 8.3, df 35.6, $P=.007$), with bursts clearly below and fizzles right at our critical threshold.